OpenStudy (anonymous):

a manufacturer of bird feeders cuts spacers from a tube that is 6 6/11 inches long. How many spacers can be cut from the tube if each spacer must be 6/11

OpenStudy (anonymous):

i think you are being asked for\[6\tfrac{6}{11}\div \frac{6}{11}\]

OpenStudy (anonymous):

I think its 6 spacers

OpenStudy (anonymous):

one way to do this is to first write \[6\tfrac{6}{11}\] as \[\frac{72}{11}\] then \[\frac{72}{11}\div \frac{6}{11}=\frac{72}{11}\times \frac{11}{6}\] cancel away

OpenStudy (anonymous):

i think you guessed right?